40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 4394     Entry time: Thu Mar 10 01:28:47 2011
Author: joe, jamie, rana, chris 
Type: Summary 
Category: CDS 
Subject: SimSuspension ! 

Today was a banner day for Simulated Plants.

Joe and Jamie have been working to get it all happening and this afternoon we started stuffing filters into the plant to make it act like the:

40mETMY.png

We put in the following features so far:

  1. Anti-Imaging filters (these are hacked to be approximate since the real ones are 7570 Hz LP filters and the SimAI only can have filters up to 8192 Hz).
  2. Dewhitening filters (copied from the SimDW in the SUS-ETMY screens)
  3. Coil Driver transimpedance (1 / 200 Ohms)
  4. Magnet-coil force constant (0.016 N/A)
  5. Conversion from Coil to DOF Basis
  6. All DOFs of the mechanical model are represented as simple harmonic oscillators with Q~100 and f ~ measured free swinging peaks.
  7. Signals/Noise can be injected either as force noise on the test mass or as displacement noise at the suspension point.
  8. Conversion from DOF to Shadow Sensor basis.
  9. Optical Levers (with whitening)
  10. Shadow Sensors have 2V/mm readout gain and whitening filters before being digitized by the SimADC.

We have also changed the switching logic for the SUS and SimETMs for the shadow sensor whitening. It used to be that either the hardware OR the software whitening was on. Now we have made it like all of the other whitening/antiwhitening in LIGO and the whitening/antiwhitening come on together. Joe and Jamie are going to propagate this to the other SUS. The hardware filter is a 30,100:3 (poles:zeros) whitening filter. The digital filter used to also be 30,100:3 with a DC gain = 1. I've changed the FM1 filter in the XXSEN filter banks into a 3:30 for the ETMY so that it now comes on and just compensates the hardware filter. This change should be propagated to all other SUS and the MEDM screens updated to show the new situation.

After this change, we decided to calibrate the {UL,UR,LL,LR,SD}SEN channels into units of microns. To do this we have made an FM6 filter called 'cts2um' that accounts for the OSEM gain and the ADC conversion factors. These channels are now in units of microns without applying any calibration in the DTT or Dataviewer. The plot below shows the OSEM shadow sensor time series with all damping loops ON and a very rough version of seismic noise being injected in all 6 DOFs (note that the y-axis is microns and the x-axis is seconds).

dvsim.png

Next, Jamie is adding the angular calibrations (so that SUSPIT and SUSYAW are in rads) and Chris is making vectift quality seismic noise injectors.

We also need to add coating thermal noise, suspension thermal noise, substrate thermal noise, ADC/DAC noise and a lot of MEDM screen indicators of what state we're in. I myself can't tell from the OSEM time series if its real or Sim.

redpill_bluepill.jpg

ELOG V3.1.3-